joe invested $5000 part at 11 annual interest and the rest at 13% annual interest if he receives $610 interest at the end of one year how much did he invest at each rate
Answer by fractalier(6550)

Call what he invests at 11%, x.
Then what he invests at 13% must be 5000-x.
The set up looks like this:
.11x + .13(5000-x) = 610
.11x + 650 - .13x = 610
-.02x = -40
x = $2000 at 11%
5000-x = $3000 at 13%